OpenStudy (anonymous):

X= (the square root of X + 13) + 7

jhonyy9 (jhonyy9):

x=sqrt(x+13) +7 hence correct ?

jhonyy9 (jhonyy9):

so then x-7=sqrt(x+13) square both sides (x-7)^2 = x-13 x^2 -14x +49 = x-13 x^2 - 15x +62 = 0 so this is a quadratic what you need to solve

OpenStudy (anonymous):

why does x+13 change to x-13?